# Strategic Analysis of Fair Rank-Minimizing Mechanisms with Agent Refusal Option

### v2

@@ 5 Concluding Remarks
First, we show the only-if-part. Suppose $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao^{\prime\prime}}=1$ where $R\left(o^{\prime\prime},r_{a}\right)<R\left(\varnothing,r_{a}\right)$. Toward a contradiction, suppose $y\notin\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$. We first consider the case where there is $y^{\prime}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$ that satisfies $y_{ao^{\prime}}^{\prime}=1$ for $o^{\prime}$ such that $R\left(o^{\prime},r_{a}\right)<R\left(\varnothing,r_{a}\right)$. By the construction of $d_{a}$,
@@
We second consider the case where there is no such a deterministic assignment. Since $\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)\neq\emptyset$ (by Corollary 1), there is $y^{\prime}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$ such that $y_{ao^{\prime}}^{\prime}=1$ for $o^{\prime\prime}$ such that $R\left(o^{\prime\prime},r_{a}\right)\geq R\left(\varnothing,r_{a}\right)$. In this case,
@@
Now, we consider $y^{\prime\prime}$ such that $y_{a\varnothing}^{\prime\prime}=1,$ $y_{ao}^{\prime\prime}=0$ for all $o\in\mathcal{O}\setminus\left\{\varnothing\right\}$ and $y_{a^{\prime}o}^{\prime\prime}=y_{a^{\prime}o}$ for all $a^{\prime}\in\mathcal{A}\setminus\left\{a\right\}$ and $o\in\mathcal{O}$. Since $q_{\varnothing}\geq\left|\mathcal{A}\right|$, $y^{\prime\prime}$ is an assignment. Moreover, since $R\left(o^{\prime\prime},r_{a}\right)\geq R\left(\varnothing,r_{a}\right)$,
@@
$RV\left(y^{\ast},\left(r_{a},p_{-a}\right)\right)>RV\left(y^{\prime\prime},\left(r_{a},p_{-a}\right)\right)$
contradicts $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$.
Second, we show the if-part. Suppose $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$. We show $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$. Suppose not; that is, there is $y^{\prime}\in\mathcal{Y}$ such that $RV\left(y^{\prime},\left(r_{a},p_{-a}\right)\right)<RV\left(y,\left(r_{a},p_{-a}\right)\right)$. Then, $y_{a\bar{o}}^{\prime}=1$ for $R\left(\bar{o},r_{a}\right)\leq R\left(\varnothing,r_{a}\right)$.
First, we assume $R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)$. Then, $RV\left(y^{\prime},\left(r_{a},p_{-a}\right)\right)=RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right)$. However,
$RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right)=RV\left(y^{\prime},\left(r_{a},p_{-a}\right)\right)<RV\left(y,\left(r_{a},p_{-a}\right)\right)=RV\left(y,\left(d_{a},p_{-a}\right)\right),$
contradicting $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
Second, we assume $\bar{o}=\varnothing$. Then, since
$\mathop{\textstyle\sum}\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime}}\geq\left|\mathcal{A}\right|,$ (11)
there is $\hat{o}$ such that $R\left(\hat{o},d_{a}\right)\leq R\left(o^{\prime},d_{a}\right)$. Thus, for $y^{\prime\prime}\in\mathcal{Y}$ such that $y_{a\hat{o}}^{\prime\prime}=1$ and $y_{a^{\prime}\tilde{o}}^{\prime\prime}=y_{a^{\prime}\tilde{o}}^{\prime}$ for all $a\in\mathcal{A}$ and all $\tilde{o}\in\mathcal{O}\setminus\left\{\hat{o}\right\}$, $RV\left(y^{\prime\prime},\left(d_{a},p_{-a}\right)\right)\leq RV\left(y^{\prime},\left(r_{a},p_{-a}\right)\right)$. However,
$RV\left(y^{\prime\prime},\left(d_{a},p_{-a}\right)\right)\leq RV\left(y^{\prime},\left(r_{a},p_{-a}\right)\right)<RV\left(y,\left(r_{a},p_{-a}\right)\right)=RV\left(y,\left(d_{a},p_{-a}\right)\right),$
contradicting $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
Therefore, $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$, then $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$. Q.E.D.
@@
Suppose that $y^{\ast}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao}^{\ast}=1$ with $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$. Then, $\left|\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)\right|\leq\left|\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)\right|$.
Proof. First, we show that if $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$, then either $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ or $y^{\prime}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right),$ such that $y_{a\varnothing}^{\prime}=1$ and $y_{a^{\prime}o}^{\prime}=y_{a^{\prime}o}$ for all $a^{\prime}\in\mathcal{A}\setminus\left\{a\right\}$ and $o\in\mathcal{O}$.
@@
$RV\left(y,\left(d_{a},p_{-a}\right)\right)\leq RV\left(y^{\ast},\left(d_{a},p_{-a}\right)\right)=RV\left(y^{\ast},\left(r_{a},p_{-a}\right)\right)<RV\left(y,\left(r_{a},p_{-a}\right)\right).$
Then, by the construction of $d_{a}$, $y_{ao^{\prime}}=1\,$such that $R\left(o^{\prime},r_{a}\right)\geq R\left(\varnothing,r_{a}\right)$. Since $q_{\varnothing}\geq\left|\mathcal{A}\right|$, $y^{\prime}$ is an assignment such that $y_{a\varnothing}^{\prime}=1,$ $y_{ao}^{\prime}=0$ for all $o\in\mathcal{O}\setminus\left\{\varnothing\right\}$ and $y_{a^{\prime}o}^{\prime}=y_{a^{\prime}o}$ for all $a^{\prime}\in\mathcal{A}\setminus\left\{a\right\}$ and $o\in\mathcal{O}$. Since $R\left(o^{\prime},d_{a}\right)\geq R\left(\varnothing,r_{a}\right)$ (by the construction of $d_{a}$), we have
@@
Finally, we assume $\left|\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)\right|\geq 2$ and arbitrarily let $y_{1},y_{2}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$ such that $y_{1}\neq y_{2}$. For $i=1,2$, let $y_{i}^{\prime}$ be such that $\left(y_{i}^{\prime}\right)_{a\varnothing}=1$ and $\left(y_{i}^{\prime}\right)_{a^{\prime}o}=\left(y_{i}\right)_{a^{\prime}o}$ for all $a^{\prime}\in\mathcal{A}\setminus\left\{a\right\}$ and $o\in\mathcal{O}$. We show $y_{1}^{\prime}\neq y_{2}^{\prime}$. Suppose not; that is, $y_{1}^{\prime}=y_{2}^{\prime}$. Then, $\left(y_{1}\right)_{a^{\prime}o}=\left(y_{2}\right)_{a^{\prime}o}$ for all $a^{\prime}\in\mathcal{A}\setminus\left\{a\right\}$ and $o\in\mathcal{O}$. Since $y_{1}\neq y_{2}$, we have $o_{1}\neq o_{2}$ where $\left(y_{1}\right)_{ao_{1}}=1$ and $\left(y_{2}\right)_{ao_{2}}=1$. Without loss of generality, we assume $R\left(o_{1},d_{a}\right)<R\left(o_{2},d_{a}\right)$. However, $R\left(o_{1},d_{a}\right)<R\left(o_{2},d_{a}\right)$ and $\left(y_{1}\right)_{a^{\prime}o}=\left(y_{2}\right)_{a^{\prime}o}$ for all $a^{\prime}\in\mathcal{A}\setminus\left\{a\right\}$ and $o\in\mathcal{O}$ imply
@@
By these results, if $\left\{y_{1},y_{2},\cdots,y_{n}\right\}=\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right),$ then $\left\{y_{1}^{\prime\prime},y_{2}^{\prime\prime},\cdots,y_{n}^{\prime\prime}\right\}\subseteq\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ where $y_{i}^{\prime\prime}$ is either $y_{i}$ or $y_{i}^{\prime}$ constructed above, because $y_{1}^{\prime},y_{2}^{\prime},\cdots,y_{n}^{\prime}$ are distinct. Q.E.D.
@@
$\left|\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)\right|\geq\left|\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)\right|\text{.}$
Moreover, any $y^{\ast}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao}^{\ast}=1$ with $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$ is also an element of $\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
@@
Fix $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ arbitrarily. Let $r_{a}\in\mathcal{P}$ and $o,o^{\prime}\in\mathcal{O}$ satisfy the condition in Proposition 3. Moreover, we let $d_{a}\in\mathcal{D}_{a}$ be such that $R\left(\varnothing,r_{a}\right)=R\left(o^{\prime},d_{a}\right)$.
@@
Let $y$ and $y^{\prime}$ be such that $y_{a\varnothing}=y_{ao^{\prime}}^{\prime}=1$ and $y_{a^{\prime}\bar{o}}=y_{a^{\prime}\bar{o}}^{\prime}$ for all $a\in\mathcal{A}$ and all $\bar{o}\in\mathcal{O}\setminus\left\{\varnothing,o^{\prime}\right\}$. Then, $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ if and only if $y^{\prime}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
@@
First, suppose $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$. Suppose not; that is, there is $y^{\prime\prime}\in\mathcal{Y}$ such that $RV\left(y^{\prime\prime},\left(d_{a},p_{-a}\right)\right)<RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right)$. Then, since (11) and Lemma 2 are satisfied, $o\in\mathcal{O}$ such that $y_{ao}^{\prime\prime}=1$ and $R\left(o,d_{a}\right)\leq R\left(o^{\prime},d_{a}\right)$. Let
$o^{\prime\prime\prime}=o\text{ if }o\neq o^{\prime},\text{ }o^{\prime\prime\prime}=\varnothing\text{ if }o=o^{\prime}\text{.}$
Moreover, let $y^{\prime\prime\prime}$ be such that $y_{ao^{\prime\prime\prime}}^{\prime\prime\prime}=1$ and $y_{a^{\prime}o}^{\prime\prime\prime}=y_{a^{\prime}o}^{\prime\prime}$ for any other pairs $(a^{\prime},o)\neq\left(a,o^{\prime\prime\prime}\right)$. Then,
@@
Second, suppose $y^{\prime}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$. Suppose not; that is, there is $y^{\prime\prime}\in\mathcal{Y}$ such that $RV\left(y^{\prime\prime},\left(r_{a},p_{-a}\right)\right)<RV\left(y,\left(r_{a},p_{-a}\right)\right)$. Then, since (11) and Lemma 2 are satisfied, $o\in\mathcal{O}$ such that $y_{ao}^{\prime\prime}=1$ and $R\left(o,r_{a}\right)\leq R\left(\varnothing,r_{a}\right)$. Let
$o^{\prime\prime\prime}=o\text{ if }o\neq\varnothing,\text{ }o^{\prime\prime\prime}=\bar{o}\text{ if }o=\varnothing\text{,}$
where $\bar{o}$ is such that $\sum y_{a\bar{o}}^{\prime\prime}<q_{\bar{o}}$ and $R\left(\bar{o},d_{a}\right)\leq R\left(o^{\prime},d_{a}\right)$. There exist such an $\bar{o}$, because (11) is satisfied. Then, let $y^{\prime\prime\prime}$ be such that $y_{ao^{\prime\prime\prime}}^{\prime\prime\prime}=1$ and $y_{a^{\prime}o}^{\prime\prime\prime}=y_{a^{\prime}o}^{\prime\prime}$ for any other pairs $(a^{\prime},o)\neq\left(a,o^{\prime\prime\prime}\right)$. Then,
@@
First, suppose that $p$ satisfies all of Conditions (i), (ii) and (iii). Then, for any two agents $\hat{a}$ and $\bar{a}$ belonging to different groups, $p_{\hat{a}}$ and $p_{\bar{a}}$ are not essentially the same. Moreover, for any two agents $\hat{a}$ and $\bar{a}$ belonging to the second group or third group, $\left(f^{M}\left(p\right)\right)_{\hat{a}o^{\prime}}=\left(f^{M}\left(p\right)\right)_{\bar{a}o^{\prime}}$ for all $o^{\prime}\in\mathcal{O}$.
@@
First, suppose that (3) is satisfied for that $k$ and $r_{a}^{k}\neq\varnothing$. Since (3) is satisfied for $k$, $r_{a}^{k^{\prime}}\neq\varnothing$ for all $k^{\prime}=1,\cdots,k-1$; that is, $R(\varnothing,r_{a})>k$ and $R(\varnothing,p_{a})\geq k$. In this case, since $f^{M}$ satisfies ETE (by Proposition 5), we revisit the examples in the first result of Lemma 3. When $k=1$ or $2$, trivially, we can use the example as is, because
@@
Second, suppose (3) is satisfied and $r_{a}^{k}=\varnothing$. Then, $k\geq 2$. Moreover, in this case, $p_{a}^{k^{\prime}}\neq\varnothing$ for all $k^{\prime}=1,\cdots,k$. Therefore, $R(\varnothing,p_{a})>R(\varnothing,r_{a})=k$. Let $p_{-a}$ be such that there are $q_{o}$ students represented by $a^{\prime}$ whose revealed preference is equivalent to $r_{a}(=p_{a^{\prime}})$ and the first best type of the other agents represented by $a^{\prime\prime}$ is $\varnothing$. Then, $\left(p_{a},p_{-a}\right)$ satisfies (i), (ii) and (iii) with $l=k$, because $k$ satisfies (3). Moreover, $\left(r_{a},p_{-a}\right)$ does not satisfy all of them. Therefore, in this case, $g\left(f^{M}\left(r_{a},p_{-a}\right)\right)$ is more preferable for $a$ than $g\left(f^{M}\left(p_{a},p_{-a}\right)\right)$.
Third, suppose that $k$ does not satisfy (3); that is,
$\mathop{\displaystyle\sum}\limits_{\bar{o}:R(\bar{o},p_{a})<k}q_{\bar{o}}=\mathop{\displaystyle\sum}\limits_{\bar{o}:R(\bar{o},r_{a})<k}q_{\bar{o}}\geq\left|\mathcal{A}\right|$
In this case, $r_{a}$ and $p_{a}$ are essentially the same. Since $f^{M}$ is a fair rank-minimizing mechanism (by Proposition 5),

### v3

@@ 5 Concluding Remarks
Suppose $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao}=1$ where $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$. Toward a contradiction, suppose $y\notin\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
We first consider the case where there is $y^{\prime}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$ that satisfies $y_{ao^{\prime}}^{\prime}=1$ for $o^{\prime}$ such that $R\left(o^{\prime},r_{a}\right)<R\left(\varnothing,r_{a}\right)$. By the construction of $d_{a}$,
@@
We second consider the case where there is no such a deterministic assignment. Since $\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)\neq\emptyset$ (by Corollary 1), there is $y^{\prime}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$ such that $y_{ao^{\prime}}^{\prime}=1$ for $o$ such that $R\left(o,r_{a}\right)\geq R\left(\varnothing,r_{a}\right)$. In this case,
@@
Now, we consider $y^{\prime\prime}$ such that $y_{a\varnothing}^{\prime\prime}=1,$ $y_{a\bar{o}}^{\prime\prime}=0$ for all $\bar{o}\in\mathcal{O}\setminus\left\{\varnothing\right\}$ and $y_{a^{\prime}\bar{o}}^{\prime\prime}=y_{a^{\prime}\bar{o}}$ for all $a^{\prime}\in\mathcal{A}\setminus\left\{a\right\}$ and all $\bar{o}\in\mathcal{O}$. Since $q_{\varnothing}\geq\left|\mathcal{A}\right|$, $y^{\prime\prime}$ is an assignment. Moreover, since $R\left(o,r_{a}\right)\geq R\left(\varnothing,r_{a}\right)$,
@@
$RV\left(y,\left(r_{a},p_{-a}\right)\right)>RV\left(y^{\prime\prime},\left(r_{a},p_{-a}\right)\right)$
contradicts $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$. Q.E.D.
@@
Suppose that there is $y^{\ast}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao}^{\ast}=1$ and $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$. Then, $\left|\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)\right|\leq\left|\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)\right|$.
Proof. Suppose that there is $y^{\ast}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao}^{\ast}=1$ and $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$. Then, by Claim 1, $y^{\ast}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
Next, we show that if $y\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$, then either $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ or $y^{\prime}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right),$ such that $y_{a\varnothing}^{\prime}=1$ and $y_{a^{\prime}o}^{\prime}=y_{a^{\prime}o}$ for all $a^{\prime}\in\mathcal{A}\setminus\left\{a\right\}$ and $o\in\mathcal{O}$. Since $q_{\varnothing}\geq\left|\mathcal{A}\right|$, $y^{\prime}\in\mathcal{Y}$.
@@
$RV\left(y,\left(d_{a},p_{-a}\right)\right)=RV\left(y^{\ast},\left(d_{a},p_{-a}\right)\right)=RV\left(y^{\ast},\left(r_{a},p_{-a}\right)\right)<RV\left(y,\left(r_{a},p_{-a}\right)\right).$
Then, by the construction of $d_{a}$, $y_{ao^{\prime}}=1\,$such that $R\left(o^{\prime},r_{a}\right)\geq R\left(\varnothing,r_{a}\right)$ and $R\left(o^{\prime},d_{a}\right)\geq R\left(\varnothing,r_{a}\right)$. Then, we have
@@
Finally, we assume $\left|\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)\right|\geq 2$ and arbitrarily let $y_{1},y_{2}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$ such that $y_{1}\neq y_{2}$. For $i=1,2$, let $y_{i}^{\prime}$ be such that $\left(y_{i}^{\prime}\right)_{a\varnothing}=1$ and $\left(y_{i}^{\prime}\right)_{a^{\prime}o}=\left(y_{i}\right)_{a^{\prime}o}$ for all $a^{\prime}\in\mathcal{A}\setminus\left\{a\right\}$ and $o\in\mathcal{O}$.
We show $y_{1}^{\prime}\neq y_{2}^{\prime}$. Suppose not; that is, $y_{1}^{\prime}=y_{2}^{\prime}$. Then, $\left(y_{1}\right)_{a^{\prime}o}=\left(y_{2}\right)_{a^{\prime}o}$ for all $a^{\prime}\in\mathcal{A}\setminus\left\{a\right\}$ and $o\in\mathcal{O}$. Since $y_{1}\neq y_{2}$, we have $o_{1}\neq o_{2}$ where $\left(y_{1}\right)_{ao_{1}}=1$ and $\left(y_{2}\right)_{ao_{2}}=1$. Without loss of generality, we assume $R\left(o_{1},d_{a}\right)<R\left(o_{2},d_{a}\right)$. However, $R\left(o_{1},d_{a}\right)<R\left(o_{2},d_{a}\right)$ and $\left(y_{1}\right)_{a^{\prime}o}=\left(y_{2}\right)_{a^{\prime}o}$ for all $a^{\prime}\in\mathcal{A}\setminus\left\{a\right\}$ and $o\in\mathcal{O}$ imply
@@
By these results, if $\left\{y_{1},y_{2},\cdots,y_{n}\right\}=\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right),$ then $\left\{y_{1}^{\prime\prime},y_{2}^{\prime\prime},\cdots,y_{n}^{\prime\prime}\right\}\subseteq\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ where $y_{i}^{\prime\prime}$ is either $y_{i}$ or $y_{i}^{\prime}$ constructed above. Since $y_{1}^{\prime},y_{2}^{\prime},\cdots,y_{n}^{\prime}$ are distinct, $y_{1}^{\prime\prime},y_{2}^{\prime\prime},\cdots,y_{n}^{\prime\prime}$ are also distinct and thus $\left|\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)\right|\leq\left|\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)\right|$. Q.E.D.
@@
$\left|\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)\right|\leq\left|\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)\right|\text{.}$
Moreover, by Claim 1, any $y^{\ast}\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ such that $y_{ao}^{\ast}=1$ with $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$ is also an element of $\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
@@
$R\left(o,r_{a}\right)$ $<$ $R\left(\varnothing,r_{a}\right)<R\left(o^{\prime},r_{a}\right)$
$\sum\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime}}$$\geq$ $\left|\mathcal{A}\right|,$ (9)
for some $o,o^{\prime}\in\mathcal{O}$. Fix $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ arbitrarily. Moreover, we let $d_{a}\in\mathcal{D}_{a}$ be such that $R\left(\varnothing,r_{a}\right)=R\left(o^{\prime},d_{a}\right)$.
@@
Let $y$ and $y^{\prime}$ be such that $y_{a\varnothing}=y_{ao^{\prime}}^{\prime}=1$ and $y_{a^{\prime}\bar{o}}=y_{a^{\prime}\bar{o}}^{\prime}$ for all $a^{\prime}\in\mathcal{A}$ and all $\bar{o}\in\mathcal{O}$. Then, $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$ if and only if $y^{\prime}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$.
@@
First, suppose $y\in\mathcal{Y}^{\ast}\left(r_{a},p_{-a}\right)$. Then, by (9), $y^{\prime}$ is an allocation. Suppose not; that is, there is $y^{\prime\prime}\in\mathcal{Y}$ such that $RV\left(y^{\prime\prime},\left(d_{a},p_{-a}\right)\right)<RV\left(y^{\prime},\left(d_{a},p_{-a}\right)\right)$. Then, by (9), $o^{\prime\prime}$ such that $y_{ao^{\prime\prime}}^{\prime\prime}=1$ satisfies $R\left(o,d_{a}\right)\leq R\left(o^{\prime},d_{a}\right)$. Let
$o^{\prime\prime\prime}=o^{\prime\prime}\text{ if }o\neq o^{\prime},\text{ }o^{\prime\prime\prime}=\varnothing\text{ if }o^{\prime\prime}=o^{\prime}\text{.}$
Moreover, let $y^{\prime\prime\prime}$ be such that $y_{ao^{\prime\prime\prime}}^{\prime\prime\prime}=1$ and $y_{a^{\prime}\bar{o}}^{\prime\prime\prime}=y_{a^{\prime}\bar{o}}^{\prime\prime}$ for all $a^{\prime}\in\mathcal{A}$ and all $\bar{o}\in\mathcal{O}$. Then,
@@
Second, suppose $y^{\prime}\in\mathcal{Y}^{\ast}\left(d_{a},p_{-a}\right)$. Suppose not; that is, there is $y^{\prime\prime}\in\mathcal{Y}$ such that $RV\left(y^{\prime\prime},\left(r_{a},p_{-a}\right)\right)<RV\left(y,\left(r_{a},p_{-a}\right)\right)$. Then, by (9), $o^{\prime\prime}\in\mathcal{O}$ such that $y_{ao^{\prime\prime}}^{\prime\prime}=1$ and $R\left(o^{\prime\prime},r_{a}\right)\leq R\left(\varnothing,r_{a}\right)$. Let
$o^{\prime\prime\prime}=o^{\prime\prime}\text{ if }o^{\prime\prime}\neq\varnothing,\text{ }o^{\prime\prime\prime}=\bar{o}\text{ if }o^{\prime\prime}=\varnothing\text{,}$
where $\bar{o}$ satisfies $\sum y_{a\bar{o}}^{\prime\prime}<q_{\bar{o}}$ and $R\left(\bar{o},d_{a}\right)\leq R\left(o^{\prime},d_{a}\right)$. There exist such an $\bar{o}$, because (9) is satisfied. Then, let $y^{\prime\prime\prime}$ be such that $y_{ao^{\prime\prime\prime}}^{\prime\prime\prime}=1$ and $y_{a^{\prime}\bar{o}}^{\prime\prime\prime}=y_{a^{\prime}\bar{o}}^{\prime\prime}$ for all $a^{\prime}\in\mathcal{A}$ and all $\bar{o}\in\mathcal{O}$. Then,
@@
First, suppose that $p$ satisfies all of Conditions (i), (ii) and (iii). Then, for any two agents $a_{1}$ and $a_{2}$ belonging to different groups, $p_{a_{1}}$ and $p_{a_{2}}$ are not essentially the same. Moreover, for any two agents $a_{1}$ and $a_{2}$ belonging to the same group (either the second group or the third group), $\left(f^{M}\left(p\right)\right)_{a_{1}o^{\prime}}=\left(f^{M}\left(p\right)\right)_{a_{2}o^{\prime}}$ for all $o^{\prime}\in\mathcal{O}$.
@@
First, suppose $k\leq\bar{k}\left(p_{a}\right)$ and $r_{a}^{k}\neq\varnothing$. Since $k\leq\bar{k}\left(p_{a}\right)$, $r_{a}^{k^{\prime}}\neq\varnothing$ for all $k^{\prime}=1,\cdots,k-1$; that is, $R(\varnothing,r_{a})>k$ and $R(\varnothing,p_{a})\geq k$. In this case, since $f^{M}$ satisfies ETE (by Proposition 5), we revisit the examples in the first result of Lemma 3. When $k=1$ or $2$, trivially, we can use the example as is, because
@@
Second, suppose $k\leq\bar{k}\left(p_{a}\right)$ and $r_{a}^{k}=\varnothing$. Then, since we assume $r_{a}^{1}\neq\varnothing$, $k\geq 2$. Moreover, in this case, $p_{a}^{k^{\prime}}\neq\varnothing$ for all $k^{\prime}=1,\cdots,k$. Therefore, $R(\varnothing,p_{a})>R(\varnothing,r_{a})=k$. Let $p_{-a}$ be such that there are $q_{o}$ students represented by $a^{\prime}$ whose revealed preference is equivalent to $r_{a}(=p_{a^{\prime}})$ and the first best type of the other agents represented by $a^{\prime\prime}$ is $\varnothing$. Then, $\left(p_{a},p_{-a}\right)$ satisfies (i), (ii) and (iii) with $l=k$, because $k\leq\bar{k}\left(p_{a}\right)$. Moreover, $\left(r_{a},p_{-a}\right)$ does not satisfy all of them. Therefore, in this case, $g\left(f^{M}\left(r_{a},p_{-a}\right)\right)$ is more preferable for $a$ than $g\left(f^{M}\left(p_{a},p_{-a}\right)\right)$.
Third, suppose $k>\bar{k}\left(p_{a}\right)$. In this case, $r_{a}$ and $p_{a}$ are essentially the same. Since $f^{M}$ satisfies ETE (by Proposition 5),